OpenStudy (anonymous):

\[\int \frac{1}{(1-u^{4})^{\frac{1}{4}}} du \]

OpenStudy (jamesj):

This can't be calculated as an elementary function, but is something rather complicated.

OpenStudy (anonymous):

Can you tell me the integration of \[\int \sqrt{tan(x)} dx\] ??

OpenStudy (jamesj):

Let u = sqrt(tan x) and then create a rational function. You then need to use partial fractions and grind away at it a bit.

OpenStudy (anonymous):

then what do?? \[\int \frac{2u^{2}}{1+u^{4}}\]

OpenStudy (anonymous):

and how to solve first one

OpenStudy (jamesj):

With respect to the integral you've just written down in u, you now need to be rather inventive with the denominator of 1 + u^4 and write it as \[(u^2 + 1)^2 - (\sqrt{2}x)^2\] Now apply difference of squares and a partial fraction decomposition. With respect to the first integral, I don't have anything concrete for you other than the Wolfram solution. Even differentiating that answer to confirm it is the integral would be gruesome.

OpenStudy (jamesj):

correction: all in u obviously \[(u^2 + 1)^2 - (\sqrt{2}u)^2\]